OpenStudy (jhannybean):

Can someone explain to me how Electron affinity and ionization energy works?

OpenStudy (anonymous):

ionization energ is the energy required to remove an electron from atom or ion in the gas phase. Electron affinity is need when an atom in the gas phase acquires an electron.

OpenStudy (anonymous):

does it help? or I did not answer the question

OpenStudy (jhannybean):

So for example: \[\sf Ca~ (g) \rightarrow Ca^+ (g) ~ + ~ e^{-}\] This is Ionization energy, right? The `+` indicates we have taken away an electron? To show that we write it as \(\sf +~e^{-}\)?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Awesome :) Then the reverse (electron affinity) would work as ...\[\sf Ca^+ (g) +~ e^- \rightarrow Ca~(g) \]

OpenStudy (anonymous):

yep

OpenStudy (jhannybean):

Just reversing it I'm guessing... and this is hypothetical :D

OpenStudy (jhannybean):

Awesome!

OpenStudy (anonymous):

good night

OpenStudy (jhannybean):

I've got a test today on enthalpy changes, etc... trying to understand all the types of enthalpy atm

OpenStudy (jhannybean):

Goodnight!

OpenStudy (jhannybean):

Oh! May Iask one last question?

OpenStudy (jhannybean):

If you're looking for bond energy... and the formation gives you half of that amount, your total reaction would be multiplied by how many MOLES of the atom there actually are, right?

OpenStudy (anonymous):

Give me example

OpenStudy (jhannybean):

I'll pull out my example here.. I'm looking for the bond energy of \(\sf O_2\), but my formation equation is \[\Delta H = 2Na(s)~ +~ \frac{1}{2}O_2 (g) \rightarrow ~ Na_2O(s)\]

OpenStudy (jhannybean):

This is not all the parts to it, but according to this, the energy I would find here would only be half the amount.. so I would have to work backwards to realize that \[\sf O_2 \rightarrow 2O\] right?

OpenStudy (anonymous):

1/2 mean half mole of the oxygen molecule. Oxygen is diatomic molecule that is why it write O2

OpenStudy (anonymous):

Half of the mole has nothing to do with half energy.

OpenStudy (anonymous):

it just balance equation

OpenStudy (jhannybean):

\[\sf Na(s) \rightarrow Na(g)~~~\Delta H = 107\]\[\sf Na(g) \rightarrow Na^+ +e^-~~~\Delta H =496 \]\[\sf O(g) +e^- \rightarrow O^{-}(g) ~~ \Delta H = -141\]\[\sf O^- +e^- \rightarrow O^{2-}(g) ~~ \Delta H = 878\]\[\sf 2Na(s) + \frac{1}{2}O_2(g) \rightarrow Na_2O ~~ \Delta H = -416\]\[\sf 2Na^+ O^{2-} \rightarrow Na_2O(s) ~~ \Delta H = -2608\]

OpenStudy (jhannybean):

And here i am told to find bond energy of \(O_2\)

OpenStudy (jhannybean):

(i forgot some of the phases, quickly wrote it out)

OpenStudy (jhannybean):

So \(\sf \Delta H = 107 +496 + (-141) +878 + (-2608) + \Delta H_{BE} = -416\)?

OpenStudy (jhannybean):

The answer choices were: a. 356 b. 498 c. 426 d. 249 e. 852

OpenStudy (jhannybean):

When I solved for this, I got 249 , but the answer was double of it, 498

OpenStudy (jhannybean):

I wanted to know why it was doubled..

OpenStudy (anonymous):

NaO2. Sodium requires 2 oxygen that is why doubled

OpenStudy (anonymous):

one oxygen is required only 249, but energy Na requires to make Nao2 is twice as much

OpenStudy (jhannybean):

Hmm...

OpenStudy (jhannybean):

Oh, if we only required half it would be something like \(NaO\) right?

OpenStudy (anonymous):

look at O-2. there are two negative ( 2 electron).

OpenStudy (jhannybean):

That's when we would get 249.

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Yeah I see that

OpenStudy (jhannybean):

I mean, two EAs.

OpenStudy (anonymous):

we are good now. I need to go to sleep

OpenStudy (jhannybean):

Oh ok, thank you very much!

OpenStudy (anonymous):

I took chemistry long time ago.....that is all I can do.

OpenStudy (jhannybean):

It was helpful :)

OpenStudy (anonymous):

I love chemistry. please come ask me in the future.

OpenStudy (jhannybean):

Willdo!

OpenStudy (anonymous):

I also can help you with math , up to pre-cal

OpenStudy (jhannybean):

:) Ok. Have a good night :D

OpenStudy (abb0t):

electron affinity is given in units of \(\sf \color{red}{\frac{J}{mol}}\) or kJ as well, in which case, some atom accepts an electron, or in otherwords, the likelihood that it will accept that electron. Ionization, as it's name implies, with "\(ion\)" in general relates to formation of cations, because you're removing the electron.